# Adaptive Testing for LLM Evaluation: A Psychometric Alternative to Static Benchmarks

### v1

@@ 4 Experiments
We evaluate the proposed ATLAS framework across five benchmarks, comparing its efficiency and accuracy against static baselines while analyzing exposure rates, overlap patterns, and the advantages of IRT-based ability estimation over raw accuracy scores.
@@ 4.1 Experimental Setup and Metrics
We compare against four static baselines that do not adapt to individual models: (1) Random sampling of 100 items uniformly from the full bank, (2) TinyBenchmarks (polo2024tinybenchmarks, ) using predetermined subsets selected via clustering without Fisher information optimization, (3) MetaBench-Primary and (4) MetaBench-Secondary (metabench, ) using curated splits that require computationally expensive iterations to identify stable subsets. Unlike these static approaches, ATLAS uses three precision thresholds ($\mathrm{SE}(\hat{\theta})\leq 0.1,0.2,0.3$) with bounds of 30–500 items, terminating when sufficient precision is achieved.
We evaluate using five metrics: (1) Mean Absolute Error (MAE) between ATLAS estimates $\hat{\theta}_{\ell}$ and full-bank references $\hat{\theta}_{\ell}^{\text{whole}}$ to assess accuracy, (2) Average Number of Items Administered to measure efficiency, (3) Item Exposure Rate quantifying how frequently each item appears across evaluations, (4) Test Overlap Rate $\bar{Q}$ measuring the average proportion of shared items between any two model evaluations, and (5) Runtime per session to assess computational efficiency. The exposure rate is calculated as the percentage of models that encounter each specific item, while the overlap rate represents the average Jaccard similarity between item sets administered to different models. Additional experimental configurations and psychometric interpretations are provided in Appendix C.
Table 1: Baseline and adaptive evaluation results (MAE and Avg. Item). Lower is better $\downarrow$.
Method WinoGrande TruthfulQA HellaSwag GSM8K ARC
MAE Items MAE Items MAE Items MAE Items MAE Items
Random 0.176 100 0.105 100 0.231 100 0.164 100 0.147 100
TinyBenchmarks 0.247 100 0.161 100 0.299 100 0.173 100 0.165 100
MetaBench-P 0.148 133 0.081 154 0.164 93 0.154 100 0.142 145
MetaBench-S 0.210 106 0.074 136 0.238 58 0.162 100 0.133 100
ATLAS0.1 0.149 78 0.067 51 0.154 39 0.159 73 0.099 77
ATLAS0.2 0.181 39 0.074 30 0.159 30 0.181 39 0.126 36
ATLAS0.3 0.183 32 0.072 30 0.161 30 0.186 32 0.128 30
Table 2: Adaptive evaluation efficiency and diversity. ATLAS maintains low item exposure rates ($10\%$) and moderate test overlap ($16-27\%$), providing natural contamination resistance while achieving fast selection times (<1<1 minute per model). Lower values indicate better performance for all metrics.
(Item #) Rate (%) Exposure Rate (%) Time (s)
WinoGrande
ATLASSE≤0.1 18.85 9.28 40.99
ATLASSE≤0.2 17.32 5.07 19.92
ATLASSE≤0.3 20.28 4.24 16.74
TruthfulQA
ATLASSE≤0.1 18.93 8.17 15.97
ATLASSE≤0.2 21.69 10.56 9.37
ATLASSE≤0.3 21.67 10.68 9.72
HellaSwag
ATLASSE≤0.1 15.76 3.90 75.52
ATLASSE≤0.2 19.28 6.23 56.93
ATLASSE≤0.3 19.25 6.85 57.06
GSM8K
ATLASSE≤0.1 22.43 8.32 45.69
ATLASSE≤0.2 22.16 5.63 24.08
ATLASSE≤0.3 26.89 8.24 19.06
ARC
ATLASSE≤0.1 19.50 10.08 30.99
ATLASSE≤0.2 19.25 5.42 13.98
ATLASSE≤0.3 21.91 9.77 11.82
4.2 Performance and Reliability Analysis
Figure 1: Comparison of subset (predicted) ability estimates against whole-bank (reference) abilities across five benchmarks. Points along the diagonal indicate perfect agreement. ATLAS maintains the closest alignment overall, particularly on TruthfulQA, ARC, HellaSwag and in the low-ability regime of GSM8K, while static baselines such as TinyBenchmarks and random show greater variance and systematic deviation.
Table 1 reveals ATLAS’s consistent efficiency advantages across all benchmarks. On HellaSwag, ATLAS achieves the best accuracy (0.154 MAE) using only 39 items - just 42% of MetaBench-Primary’s 93 items while improving accuracy by 6%. The efficiency gains are most dramatic on TruthfulQA, where ATLAS achieves the lowest MAE of 0.067 using only 33% of MetaBench-Primary’s items (51 vs 154) and outperforms both MetaBench variants. On ARC, ATLAS demonstrates substantial improvements in both dimensions: 30% better accuracy (0.099 vs 0.142 MAE) using just 53% of the items (77 vs 145). Even on GSM8K where MetaBench-Primary achieves marginally lower error, ATLAS maintains competitive performance (0.159 vs 0.154 MAE) while requiring 27% fewer items. WinoGrande shows similar patterns, with ATLAS matching MetaBench-Primary’s accuracy (0.149 vs 0.148 MAE) using only 59% of the items.
Figure 1 compares ability estimates from subsets against whole-bank references. ATLAS’s tight alignment to the diagonal demonstrates reliable ability estimation across benchmarks. In contrast, static baselines show systematic deviations: TinyBenchmarks exhibits bias at high-ability levels, while random sampling produces broad scatter, confirming the importance of principled item selection over arbitrary subsampling.
4.3 Contamination Resistance and Efficiency Benefits
Table 2 demonstrates ATLAS’s contamination resistance and efficiency advantages across multiple dimensions. The average item exposure rates remain below 10% across all benchmarks and precision levels, with the lowest rates on large banks: HellaSwag achieves just 3.9% exposure with SE $\leq 0.1$, meaning each item appears in fewer than 4% of model evaluations. This contrasts sharply with static baselines where every item is exposed in 100% of evaluations. Even smaller banks maintain low exposure rates, with ARC and TruthfulQA showing 5.4-10.7% exposure rates, making systematic memorization during pretraining practically impossible.
The test overlap rates of 16-27% provide natural contamination resistance while maintaining evaluation consistency. On HellaSwag, the 15.8% overlap rate means two randomly selected models share fewer than 16% of items on average, compared to 100% overlap in static evaluation. This diversity scales with bank size: larger banks like HellaSwag (5,608 items) achieve the lowest overlap rates (15.8-19.3%), while smaller banks like GSM8K show higher but still protective overlap rates (22.2-26.9%). The precision threshold affects this trade-off: stricter thresholds (SE $\leq 0.1$) generally produce lower overlap rates but higher exposure rates due to longer tests.
Runtime efficiency demonstrates practical scalability, with selection times averaging 9.4-75.5 seconds per model across benchmarks. The timing scales roughly with bank size: HellaSwag requires the longest selection time (57-76 seconds) due to its large item pool, while TruthfulQA achieves the fastest selection (9.4-16.0 seconds). Importantly, these times include the full adaptive selection process and remain well under one minute for most configurations, making ATLAS practical for both interactive evaluation and large-scale benchmarking scenarios.
@@ 4.3 Contamination Resistance and Efficiency Benefits
4.4 Distinguishing Low- and High-Performing Models
@@
Figure 2 demonstrates IRT’s key advantage: separating models with similar accuracy scores through ability estimates. Despite strong correlations (0.99 for GSM8K, 0.96 for HellaSwag), systematic differences emerge at performance extremes. In low-performing regimes, accuracy collapses into narrow bands (0.10-0.15) while IRT spans $\theta\approx-3$ to 1-1, distinguishing models that solve easy versus challenging items. In high-performing regimes, ceiling effects compress accuracy differences, but IRT maintains discrimination across $\theta\approx 1.5$ to $2.5$. Similar patterns are observed across TruthfulQA, WinoGrande, and ARC benchmarks, with additional score-versus-theta comparisons provided in Appendix G.
The right panels show systematic rank reordering: 23-31% of models shift >> 10 positions when ranked by IRT versus accuracy. Models performing well on hard items receive higher IRT ranks despite moderate accuracy, while those succeeding on easy items are appropriately downweighted. This enhanced discrimination provides more reliable model comparisons, especially critical in saturated performance regions where accuracy-based evaluation fails. Similar patterns of IRT superiority in distinguishing models are observed across TruthfulQA, WinoGrande, and ARC benchmarks, with additional analysis provided in Appendix G.
@@ 4.5 Items Are Not Equally Informative
Figure 3: Two models with identical accuracy (0.833) on WinoGrande receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$). Model A succeeds on harder items (darker cells on right), while Model B answers easier items (darker cells on left). IRT captures these item difficulty patterns that raw accuracy cannot.
Figure 3 demonstrates IRT’s key advantage: models with identical accuracy (0.833) receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$) based on which items they solve correctly. Under the 3PL IRT model, items differ in discrimination $a_{i}$, difficulty $b_{i}$, and guessing $c_{i}$ parameters. Model A succeeds on high-difficulty items ($b_{i}>0.5$) with strong discrimination ($a_{i}>1.5$), yielding 2.3× more Fisher information than Model B, which mainly answers easier, less discriminative items ($b_{i}<-0.5$, $a_{i}<0.8$).
IRT provides automatic quality control by weighting items according to their empirical contribution to distinguishing model abilities. In our calibrated banks, 3.2-5.7% of items exhibit negative discrimination ($a_{i}<0$), indicating systematic flaws where stronger models perform worse. For example, WinoGrande item #247 achieves 0.89 accuracy but $a_{i}=-0.43$ due to exploitable linguistic artifacts. Under raw accuracy, this flawed item contributes equally (weight = 1/N) to all scores, potentially inflating weak models. Under IRT, negative discrimination automatically down-weights its contribution by 82% (effective weight $\propto a_{i}^{2}\approx 0.18$), reducing measurement contamination and providing more reliable ability estimates than accuracy alone. Similar patterns of identical accuracy leading to different ability estimates are observed across ARC, HellaSwag, and other benchmarks, with additional heatmap visualizations provided in Appendix G.
@@ 5 Discussion
5.1 Implications for Benchmark Design
Our findings underscore several critical considerations for the design of future large language model (LLM) benchmarks. First, reduced item sets necessitate content balancing to preserve assessment validity. Content balancing entails maintaining proportional representation across distinct skill domains or cognitive competencies to ensure that the benchmark reflects the intended construct rather than overemphasizing specific subskills. Without such balance, benchmarks risk producing biased or unrepresentative evaluations. The proposed adaptive framework can be readily extended to optimize both domain coverage and measurement precision simultaneously (cheng2009maximum, ). Achieving comparable balance using static, reduced item sets would be substantially more difficult and would require extensive manual tuning or domain expertise.
Second, item quality exerts a direct influence on evaluation reliability. Our analysis reveals that 3.2–5.7% of items display negative discrimination parameters, indicating systematic flaws that undermine accuracy-based performance estimates. Although IRT automatically attenuates the influence of such items through probabilistic weighting, many existing benchmarks lack analogous quality control mechanisms. Consequently, poorly constructed or ambiguous items can distort aggregate scores and obscure genuine model differences. Future benchmark design should therefore adopt psychometric validation as standard practice, incorporating procedures such as discrimination screening and content alignment checks to enhance the interpretability and fairness of evaluation outcomes.
Third, item exposure rate and test overlap rate are essential metrics for benchmark robustness and reusability. A low exposure rate ensures that individual items are rarely reused across assessments, reducing the risk of model memorization or contamination of test content. Similarly, a low test overlap rate—the proportion of shared items between distinct benchmark forms—preserves evaluation independence and minimizes redundancy. Adaptive benchmarking intrinsically achieves lower exposure and overlap rates by tailoring item selection to each model’s estimated proficiency, generating diverse yet psychometrically comparable test forms. This adaptivity confers a distinct advantage over static benchmarks, enabling more secure, scalable, and sustainable evaluation practices in iterative model development contexts.
5.2 Limitations and Future Work
Despite notable efficiency gains, our framework has several limitations. Initial calibration relies on a representative model population, which may become outdated as architectures evolve. Nonetheless, adaptive testing supports incremental updates: new model responses can be incorporated to refine item parameters and maintain calibration accuracy over time. The current implementation remains limited to multiple-choice formats, while generative or open-ended tasks require alternative scoring and modeling strategies. Moreover, our framework is unidimensional, assuming a single latent proficiency dimension across all items, whereas LLM capabilities are inherently multidimensional. Future work should extend the framework to multidimensional IRT formulations that jointly model reasoning, factuality, and linguistic ability.
To facilitate broader adoption, we have implemented a modular scoring interface that allows users to define custom evaluation functions (e.g., determining whether a model answers a selected item correctly). Future work will explore online or Bayesian calibration techniques for continuous item updating, multidimensional modeling to capture diverse LLM capabilities, and hybrid adaptive–generative evaluation for open-ended tasks. We also plan to extend the system with interfaces for cross-model benchmarking, enhanced item exposure control, and visualization tools that improve interpretability and diagnostic insight.
@@ 6 Conclusion
Adaptive testing replaces exhaustive benchmarking with efficient, contamination-resistant measurement. Our ATLAS framework reduces items by 90% while maintaining precision, separates models with identical accuracies using IRT-based ability estimates, and limits contamination through << 10% item exposure rates. With benchmark saturation and contamination risks increasing, adaptive testing provides a rigorous alternative, applying measurement science principles instead of static averaging.
@@ Appendix B Comparison of IRT-Based Benchmark Methods
Table 3: Comparison of IRT-Based Benchmark Methods for LLMs
@@
Efficiency Fixed (100 items) Fixed (100 items) Fewer items, comparable accuracy
@@
Appendix C Detailed Experimental Setup and Metrics
C.1 Benchmarks and Datasets

### v2

@@ 4 Experiments
We evaluate the proposed ATLAS framework across five benchmarks, comparing its efficiency and accuracy against static baselines such as random sampling, TinyBenchmarks (Polo et al., 2024), MetaBench-Primary and MetaBench-Secondary (Kipnis et al., 2025). We report performance-based and efficiency-based metrics, with full metric definitions provided in Appendix E.
@@ 4.1 Experimental Setup and Metrics
We compare against four static baselines that do not adapt to individual models: (1) Random sampling of 100 items uniformly from the full bank, (2) TinyBenchmarks (Polo et al., 2024) using predetermined subsets selected via clustering without Fisher information optimization, (3) MetaBench-Primary and (4) MetaBench-Secondary (Kipnis et al., 2025) using curated splits that require computationally expensive iterations to identify stable subsets. Unlike these static approaches, ATLAS uses three precision thresholds ($\mathrm{SE}(\hat{\theta})\leq 0.1,0.2,0.3$) and an item bound of 30–500, terminating adaptively when the required precision is achieved or the maximum test length is reached. Additional experimental configurations are provided in Appendix D.
We evaluate each method primarily in ability space, comparing ATLAS-estimated abilities $\hat{\theta}_{\ell}$ with full-bank abilities $\hat{\theta}^{\text{whole}}_{\ell}$ (See Table 2). We report four metrics: (1) Mean Absolute Error (MAE) to measure estimation accuracy; (2) Standard Error (SE) of the absolute errors across models to quantify stability; (3) Average Test Length, the number of items administered per model; and (4) Information Efficiency Score (IES), which jointly reflects accuracy and item usage relative to a 100-item random baseline (values <1<1 indicate higher efficiency). More details of these metrics are provided in Appendix E. We also provide accuracy-space evaluations, comparing reconstructed accuracies to full-bank raw accuracies (See Table 3). Additional evaluation metrics and their results, including Item Exposure Rate, Test Overlap Rate, and Selection Time are provided in Appendix G.2.
@@
Figure 3: Two models with identical accuracy (0.833) on WinoGrande receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$). Difficulty levels are ordered from easiest (D1) to hardest (D10). Model A performs better on harder items (darker cells on right), while Model B primarily succeeds on easier ones (darker cells on left), highlighting how IRT distinguishes latent ability differences that accuracy fails to capture.
4.2 Performance and Reliability Analysis
Table 2 presents a comparison of whole-bank ability estimates and subset-based estimates across benchmarks (results for GSM8K and ARC are provided in Appendix G.1). ATLAS consistently delivers the strongest accuracy–efficiency tradeoff among all methods. Across every benchmark, an ATLAS variant achieves the lowest Information Efficiency Score (IES), indicating that it achieves the best accuracy-efficiency trade-off. For example, ATLAS attains the best MAE on TruthfulQA (0.064 with 48 items) and HellaSwag (0.157 with 41 items), and matches the performance of MetaBench-Primary on WinoGrande while requiring nearly half as many items (70 vs. 133). Even in more challenging settings such as GSM8K and ARC (Appendix G.1), ATLAS maintains low MAE with item counts as small as 30–36, outperforming all static baselines in information efficiency. In contrast, static subsets such as TinyBenchmarks and MetaBench exhibit inconsistent performance. They are strong on some datasets but poor on others, with substantially higher IES values. Overall, these results demonstrate that adaptive item selection enables ATLAS to deliver high-fidelity ability estimates while dramatically reducing evaluation cost, achieving reliable performance that fixed subsets fail to match.
Figure 1 plots subset-based ability estimates against whole-bank references across all five benchmarks. Across datasets, ATLAS shows the closest alignment to this identity line, with tightly clustered points and minimal systematic drift, reflecting stable and high-fidelity ability estimation. In contrast, static baselines exhibit benchmark-dependent failures. Random sampling yields the widest scatter on every benchmark. TinyBenchmarks consistently deviate from the identity line in the extreme high or low-ability regime, especially on ARC and TruthfulQA. MetaBench-Primary and MetaBench-Secondary performs reasonably on GSM8K but breaks down substantially on HellaSwag, where both its primary and secondary subsets produce large downward deviations, indicating poor item coverage.
As shown in Table 8, ATLAS produces adaptive tests with low redundancy and high efficiency. Test overlap rates remain between 11–23%, and item exposure stays below 12% across all benchmarks, reflecting broad utilization of the item bank. End-to-end runtimes range from 9.4 to 75.5 seconds per model and scale consistently with bank size (shortest for TruthfulQA and longest for HellaSwag), confirming that ATLAS is computationally practical.
Accuracy Reconstruction.
To assess whether the estimated ability $\theta$ is consistent with traditional accuracy-based evaluation, we reconstruct each model’s expected percent correct from its estimated ability and observed responses on the reduced subset using the performance-IRT (p-IRT) estimator (Polo et al., 2024) (detailed in Appendix F). P-IRT builds on the Test Characteristic Curve (Lord and Novick, 2008; Hambleton et al., 1991) which maps a model’s ability $\hat{\theta}$ to its expected probability of correctly answering benchmark items under the calibrated 3PL model, and refines this mapping by incorporating observed responses.
As shown in Table 3, across all benchmarks, the reconstructed percent correct closely matches the raw full-bank percent correct, with MAE consistently below 5%, demonstrating that $\theta$ faithfully preserves global performance as measured by accuracy. This confirms that the latent ability estimates remain well aligned with conventional evaluation metrics. Moreover, ATLAS achieves this high reconstruction fidelity with substantially fewer administered items than competing methods. For example, on WinoGrande, ATLAS0.1 attains the lowest MAE (0.048) using only 70 items, compared to 133 items for MetaBench-Primary.
4.3 Discriminating Models at Performance Extremes
@@
Figure 2 illustrates IRT’s key advantage: separating models with similar accuracy via latent ability estimates. Although accuracy and ability are highly correlated (0.99 on GSM8K; 0.96 on HellaSwag), systematic differences appear at performance extremes. In low-performing regimes, accuracy collapses into narrow bands (0.10–0.15), while IRT spans $\theta\approx-3$ to 1-1, distinguishing models that solve only easy items from those that handle harder ones. At the high end, ceiling effects compress accuracy, whereas IRT preserves discrimination across $\theta\approx 1.5$ to $2.5$.
The right panels reveal substantial rank reordering: 23–31% of models shift by more than 10 positions when ranked by IRT rather than accuracy. Models that succeed on difficult items are ranked higher, while those relying on easy items are downweighted, yielding more reliable comparisons in saturated regimes. Similar trends hold across TruthfulQA, WinoGrande, and ARC (See Appendix G).
@@ 4.4 Items Are Not Equally Informative
Figure 3 illustrates a core principle of Item Response Theory: items contribute unequally to ability estimation. Although Models A and B achieve the same raw accuracy (0.833), their response patterns differ systematically across difficulty levels, leading to substantially different latent ability estimates ($\hat{\theta}_{A}=1.2$ vs. $\hat{\theta}_{B}=0.6$). Model A performs disproportionately well on harder items (D6–D9), which are more discriminative and thus carry greater informational value for estimating high ability, whereas Model B’s correct responses are concentrated on easier items (D1–D4), which provide limited evidence of advanced competence. As a result, identical accuracy masks meaningful differences in underlying capability: two models can answer the same number of questions correctly, yet differ substantially in latent ability because correct responses to difficult items provide strong evidence of high ability, while correct responses to easy items are expected and contribute little additional information. This example highlights why accuracy alone is an insufficient evaluation metric and how IRT leverages item difficulty to recover latent ability differences that raw accuracy fail to capture. Similar patterns of identical accuracy leading to different ability estimates are observed across ARC, HellaSwag, and other benchmarks, with additional heatmap visualizations provided in Appendix G.
@@ 5 Discussion
Our results underscore several psychometric principles that should guide the design of future LLM benchmarks. First, item quality is critical for reliable evaluation. Prior work has shown that mislabeled or ambiguous items are common in existing benchmarks (Vendrow et al., 2025; Gema et al., 2025). Although IRT naturally downweights such items through probabilistic modeling, static benchmarks lack mechanisms to prevent them from being sampled. Fluid Benchmarking (Hofmann et al., 2025) highlights this issue, showing that adaptive, information-based selection dramatically reduces the exposure of low-quality items. These findings collectively reinforce the need for psychometric validation in future benchmark design, including procedures such as discrimination screening and content-alignment checks.
Second, model fit is essential for trustworthy item parameter estimation, yet it is rarely examined or reported in existing benchmark-reduction methods that apply IRT. When the underlying IRT model fits poorly, the resulting difficulty and discrimination estimates become unstable, which in turn compromises any downstream conclusions drawn from reduced item sets. Our results show that several widely used static subsets exhibit substantial misfit (Table 1), underscoring that model-fit diagnostics should be a standard requirement for any benchmark that applies IRT for item calibration or reduction. Routine reporting enables researchers to verify whether the assumed psychometric model adequately captures LLM response behavior before relying on the resulting item parameters or reduced test forms.
Third, when large item banks are partitioned for computational efficiency, linking becomes necessary to prevent scale drift. Common-person linking places items on a shared latent scale, preserving interpretability and comparability. Such principled linking strategies are crucial for scalable benchmarking and for maintaining consistency across bank updates or domains.
Finally, reduced benchmarks require explicit content balancing to maintain validity. Without constraints on domain coverage, reduced item sets may overemphasize narrow skills. Our adaptive framework naturally supports joint optimization of measurement precision and content balance (Cheng and Chang, 2009), whereas achieving similar guarantees in static subsets requires extensive manual tuning.
@@ 6 Conclusion
We presented ATLAS, an adaptive testing framework that reframes LLM evaluation by moving beyond static, accuracy-based benchmarking toward dynamic ability estimation. Through psychometric modeling and information-guided item selection, ATLAS achieves up to 90% item reduction, avoids accuracy ceiling effects, and reveals ability differences that static benchmarks overlook. Our analysis further highlights the importance of rigorous model-fit validation, item-quality assessment, and principled linking procedures for building reliable benchmarks, offering a more efficient, interpretable, and robust foundation to assess LLMs.
Impact Statement
This work aims to advance the methodology of large language model evaluation by introducing a psychometrically grounded, adaptive testing framework. By improving the reliability, efficiency, and interpretability of model assessment, our approach may support more informed benchmarking and comparison of LLMs. We do not foresee direct negative societal or ethical impacts arising from this work, as it focuses on evaluation rather than deployment or decision-making. Nonetheless, as with all benchmarking tools, care should be taken to ensure that evaluations are used responsibly and do not incentivize narrow optimization at the expense of broader model robustness.
@@ Appendix B Comparison of IRT-Based Benchmark Methods
A detailed comparison of IRT-based benchmark methods is reported in Table 4.
@@
Table 4: Comparison of IRT-Based Benchmark Methods for LLMs
Appendix C Data Preprocessing Details
C.1 Point-Biserial Correlation Formula
The point-biserial correlation (Allen and Yen, 2001) for item $i$ is defined as:
C.2 Detailed Explanation of the Common-Person Calibration Procedure
In this work, calibration refers to estimating item characteristics under the 3PL item response theory model. The purpose of calibration is to place all items on a shared difficulty scale so that performance comparisons across items and models become meaningful. Under the 3PL model, each item is described by a difficulty parameter, a discrimination parameter that captures how strongly the item differentiates between high- and low-performing models, and a guessing parameter that reflects the chance of a correct response when the model effectively guesses. When every LLM responds to the same items, their collective performance patterns allow these parameters to be estimated in a consistent way. This provides a principled way to identify which items are easy or difficult for models and which items are more or less informative. Calibrating the full benchmark at once would be computationally intensive because the 3PL model becomes more expensive to fit as the number of items grows. To make the process tractable, we divide the full item pool into several non-overlapping subsets and calibrate each subset independently. This reduces the computational load substantially, but it also means that each subset is estimated on its own internal scale. For example, the notion of “difficulty” in one subset is not automatically aligned with the notion of “difficulty” in another. A separate linking step is therefore required to place all subsets onto a shared scale.
Linking requires shared reference points known as anchors. In educational measurement, anchors are typically common items or common examinees that appear in multiple test forms. They serve as a bridge that allows independently calibrated scales to be aligned. In our setting, every LLM responds to every item in the benchmark, which means that the same population of models appears in the calibration of each item subset. The models therefore act as common persons in the traditional psychometric sense and serve as the linking anchors for the benchmark. Their relative performance across subsets provides the information needed to align the scale of each subset with the others. The linking procedure examines how the same models perform across the different subsets and adjusts each subset’s scale so that the overall performance patterns match. If a model appears stronger than its peers in one subset and shows a similar relative standing in another, then the two subsets can be placed on the same scale by aligning the average performance level and the overall spread of performance. This rescaling is then applied to the item parameters of each subset so that all items, regardless of which subset they came from, are expressed on a single, unified difficulty metric.
This form of common-person linking is particularly effective for LLM benchmarking. In human testing, it is rarely feasible for every examinee to respond to every item, and linking must rely on smaller or less reliable anchor sets. LLMs do not face constraints such as fatigue, practice effects, or time limits, which allows us to use the entire model population as a complete and stable set of anchors. This makes the linking process highly robust and enables a scalable calibration framework that achieves substantial computational efficiency while maintaining coherence across a very large item bank.
WinoGrande TruthfulQA HellaSwag GSM8K ARC
# Models used for calibration 4680 4635 3467 3775 3747
# Models used for testing 521 516 386 420 417
# Calibration subsets ($K$) 10 6 50 12 8
# Items after filtering 1045 627 5600 1306 839
Average RMSEA ($M_{2}$) 0.0565 0.0690 0.0482 0.0438 0.0595
Table 5: Statistics describing the calibration dataset, testing dataset, item-bank size after filtering, number of calibration partitions (K), and average model–data fit (RMSEA from M2) across all benchmarks.
C.3 Calibration Data, Item-Bank Partitioning, and Fit Statistics
A summary of the calibration data, item-bank partitioning, and associated fit statistics is provided in Table 5.
Appendix D Detailed Experimental Setup and Metrics
D.1 Benchmarks and Datasets